If the line through (2, 6) and (11,y2)is parallel to the line
7x − 9y = 27, find y2
I know this has something to do with slopes but i cant remember

Answer by stanbon(75887)

If the line through (2, 6) and (11,y2)is parallel to the line
7x − 9y = 27, find y2
------------------------
The slope-intercept form of the given equation is:
y = (7/9)x-3
So the slope is 7/9
---------------------------
The slope determined by the two points is:
slope = (y2-6)/(11-2) = (y2-6)/9
-------------------------------------
Equate the two to solve for y2
(y2-6)/9 = 7/9
y2-6 = 7
y2 = 13
=================
Cheers,
Stan H.